OpenStudy (anonymous):

Below are two different functions, f(x) and g(x). What can be determined about their y-intercepts? f(x) = x + 4 x g(x) −1 8 1 0 2 −4 The function f(x) has a higher y-intercept. The function g(x) has a higher y-intercept. They both have the same y-intercept. The relationship between y-intercepts cannot be determined.

OpenStudy (anonymous):

any ideas

OpenStudy (anonymous):

what's g(x) when x=0 ?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

@bridgetx516x

OpenStudy (anonymous):

It's not the third option, that I know.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

You could use the given x and y values to plot g(x) and find the y-intercept graphically.

OpenStudy (anonymous):

o

OpenStudy (anonymous):

what is the answer

OpenStudy (anonymous):

This homework is about linear equations right? if it is, then my answer would probably work. If this homework is about quadratics and and linear equations together, then , I don't think my answer would be right. anyway, if this homework is only about linear equations, then Instead of doing it graphically, you can use, say (-1,8) and (1,0) to find the gradient: (y1-y2)/(x1-x2)= gradient or m (8-0)/(-1-1)= 8/-2= -4 so -4 is the gradient. After you find the gradient you can use one point on g(x) to find the linear equation using this equation: y-y1=m(x-x1) i'll pick the point: (1,0) (you can use any coordinate on g(x)) so: y-0= -4(x-1) y= -4x + 4 so -4x+4= g(x) The y-intercept is the "+4" so f(x) and g(x) HAVE THE SAME y-intercepts after all. So I stand corrected, it IS the third option.